# Large Language Models as Attribution Regularizers for Efficient Model Training

### v1

@@ Abstract
Several challenges hinder the adoption of LLMs for tabular data tasks. Firstly, their large parameter counts demand substantial computational resources, typically reliant on GPUs, which significantly increases operational costs. Secondly, tabular learning is often employed in domains where transparency and interpretability are critical, such as healthcare and finance. In these fields, simpler and more interpretable models, such as logistic regression or decision trees, are often preferred. Although techniques exist to enhance LLM interpretability [3], they currently fall short compared to the inherent explainability of simpler models.
@@ 2.2 Large Language Models in Tabular Learning
Recently, Han et al. [21] introduced FeatLLM, a novel approach that utilizes LLMs as feature engineers. Instead of directly performing inference, FeatLLM employs code-generating LLMs to create preprocessing functions that transform the original dataset into a more suitable representation for few-shot classification. This method implements an ensemble classifier to combine insights from multiple feature transformations, improving robustness and classification accuracy. FeatLLM significantly reduces resource requirements by relying solely on pre-trained LLMs with API-level access. Moreover, FeatLLM outperforms existing fine-tuned and in-context learning approaches while maintaining lower computational costs.
@@ 3.1 Formulation
3.2 LLM Prompting and Scores Parsing
To enable Large Language Models (LLMs) to generate meaningful feature attribution scores that effectively guide downstream models during training, we developed a comprehensive prompting methodology. This approach incorporates several key components to maximize score accuracy and relevance.
@@ 3.2.1 Task and Dataset Contextualization
We first embed relevant task and dataset information within the prompt structure by incorporating key details. Task descriptions provide a concise overview of the classification objective and possible outcomes. For example: "Predict whether this patient’s breast cancer will reoccur. Yes or no?" — an approach aligned with established methodologies in recent literature [21, 23]. Feature descriptions ensure clarity by defining each dataset attribute; for instance, "Age: The age of the patient at the time of diagnosis." To handle categorical features, we apply one-hot encoding and include specific descriptions for each category. This allows the LLM to generate distinct attribution scores for individual categorical levels rather than the overall feature.
@@ 3.2.2 Score Generation Protocol
We implement a structured scoring protocol by instructing The LLM to assign an integer score between and for each feature, creating a standardized scale of feature importance. Following chain-of-thought methodologies [51], we incorporate prompts that require explicit reasoning before score assignment. This approach has demonstrated superior performance in complex reasoning tasks and improves the interpretability of the scoring process. The complete prompt template used for score generation is provided in Appendix 1
@@ 3.2.3 Score Extraction and Aggregation
To transform the LLM’s textual responses into structured numerical data we employ a secondary LLM instance configured for function calling to extract numerical scores from the primary LLM’s reasoning and output. This parser converts the semi-structured scores into a standardized list format suitable for downstream processing.
To mitigate potential variability in LLM outputs, we generate scores multiple times () and compute their mean. This ensemble approach produces the final LLM-provided feature-wise attribution vector (), enhancing the stability and reliability of the attribution signal.
This methodology creates a robust framework for leveraging LLM capabilities to generate informative feature attribution scores that can effectively guide downstream predictive models during their training process.
@@ 4.1 Few-shot Learning
We evaluate our approach on seven publicly available binary classification datasets:
@@
The datasets vary in size and complexity, and their additional basic information is outlined in Appendix 5.
@@ 4.1.2 Implementation Details
We employ three distinct foundation models as our importance score estimators: Llama 3.3 70B () [15], Gemini 2.0 Flash () [40] and GPT-4o-mini () [26]. As simple downstream models, we utilized logistic regression and a two-layer multilayer perceptron (MLP) with ReLU activation and 100 hidden units. Optimization was performed using the Adam optimizer [30] with a learning rate of and no weight decay, leveraging LAAT’s internal regularization via the factor, which was set to 100. Additionally, the number of importance score estimations for LAAT was set to 5. LAAT was executed without early stopping. Data preprocessing involved one-hot encoding categorical variables and standardizing numerical features using z-score normalization. Exact task descriptions used can be found in Appendix 6.
For traditional machine learning models (logistic regression, MLP, random forest, XGBoost, and CatBoost), we performed hyperparameter optimization through grid search with 5-fold cross-validation, with the exception of 1-shot learning scenarios where default scikit-learn [39] hyperparameters were applied. The complete baseline hyperparameter search spaces are detailed in Appendix 7. TabPFN and FeatLLM were implemented with their respective default parameter configurations as described in the original publications. Additionally, FeatLLM incorporated early stopping mechanisms in the 5-shot and 10-shot experimental settings. Due to the inherent complexity of generating multiple conditions and preprocessing functions in the FeatLLM method, Llama 3.3 exhibited inconsistencies in producing valid outputs. As a result, we excluded it from the FeatLLM experiments.
To ensure statistical robustness, we conducted ten independent experimental runs for each model.
@@ 4.1.3 Results
The results of our experiments, averaged over ten runs, are presented in Tables 1 (for non-LLM methods) and 2 (for LLM-based methods). As shown, LAAT models consistently rank among the top-performing methods, achieving 16 top scores across 21 experiments. Notably, the highest-performing variants of LAAT were and , each securing 6 top scores. Furthermore, the LAAT-trained models outperformed their non-LAAT counterparts in 15, 19, and 20 out of 21 experiments using logistic regression with Llama 3.3 70B, Gemini 2.0 Flash, and GPT-4o-mini as scoring models, respectively. For MLP, LAAT models exceeded the performance of vanilla MLP in 17, 20, and 20 experiments, respectively. These results underscore the positive impact of LLM attribution alignment on the generalization capabilities of even simple models, achieving significant improvements without the need for extensive hyperparameter tuning, which is often required for other approaches.
Dataset | Shot | LR | MLP | RF | XGB | CatBoost | TabPFN |
@@
Figure 1 presents the average ROC AUC performance across all datasets. While FeatLLM achieves the highest score in the 1-shot scenario, LAAT models follow closely, surpassing all non-LLM baselines. However, as FeatLLM’s performance plateaus beyond the initial 1-shot setting, LAAT models demonstrate a substantial advantage in both the 5-shot and 10-shot scenarios, outperforming all other methods by a significant margin. Among the scoring models evaluated, GPT-4o-mini emerged as the top performer, with Gemini 2.0 Flash-Lite following closely behind. Although Llama 3.3 70B exhibited lower performance compared to these two LLMs, it still outperformed all non-LLM baselines in the 1-shot setting, surpassed all baselines in the 5-shot setting, and maintained competitive performance in the 10-shot experiments.
@@ 4.2.1 Results
As in the previous experiment, results are presented in Tables 3 for non-LLM methods and 4 for LLM-based methods. LAAT effectively mitigates dataset biases and achieves the highest performance across all datasets, surpassing non-LLM baselines by more than 20 ROC AUC points in some cases. While FeatLLM outperforms traditional baselines, it remains significantly behind LAAT models. The superior performance of LAAT on this challenging benchmark further highlights the benefits of attribution guidance — not only in enhancing performance in data-scarce scenarios but also in addressing subtle, hard-to-detect dataset biases that could otherwise compromise the integrity of the machine learning pipeline.
@@ 4.3 Hyperparameter Analysis
We investigate the impact of varying the regularization factor and the number of importance score estimations on the final performance of LAAT models. Specifically, we assess the performance of a logistic regression model trained using LAAT alignment with importance scores provided by GPT-4o mini. The experimental results are presented in Figures 3 and 3, corresponding to the variation and the number of importance score estimations, respectively.
@@ 5 Conclusion
Beyond its standalone effectiveness, LAAT is highly versatile and can be used in conjunction with more complex models, provided they are differentiable — for example, TabPFN. Additionally, it can be combined with feature preprocessing techniques such as FeatLLM. Given the interpretability of LLM-derived importance scores, LAAT could be incorporated into interactive, chat-based interfaces, enabling human experts to refine these scores before finalizing them. This human-in-the-loop refinement could further enhance model performance by leveraging domain expertise. Furthermore, LAAT’s framework could be extended beyond tabular data to other modalities, such as images, by integrating it into concept bottleneck models [31]. This extension would allow LAAT to guide feature selection and attribution in vision tasks, improving interpretability and generalization in image-based applications. We leave the exploration of these directions to future work.
@@
We provide additional dataset information, task descriptions, score extraction prompt and hyperparameter search spaces utilized in our experiments.
Basic Dataset Information
@@
Prompt Template for Importance Score Generation
Hyperparameter Search Spaces

### v2

@@ Abstract
Several challenges hinder the adoption of LLMs for tabular data tasks. Firstly, their large parameter counts demand substantial computational resources, typically reliant on GPUs, which significantly increases operational costs. Secondly, tabular learning is often employed in domains where transparency and interpretability are critical, such as healthcare and finance. In these fields, simpler and more interpretable models, such as logistic regression or decision trees, are often preferred. Although techniques exist to enhance LLM interpretability [3, 34], they currently fall short compared to the inherent explainability of simpler models.
@@ 2.2 Large Language Models in Tabular Learning
Recently, Han et al. [22] introduced FeatLLM, a novel approach that utilizes LLMs as feature engineers. Instead of directly performing inference, FeatLLM employs code-generating LLMs to create preprocessing functions that transform the original dataset into a more suitable representation for few-shot classification. This method implements an ensemble classifier to combine insights from multiple feature transformations, improving robustness and classification accuracy. FeatLLM significantly reduces resource requirements by relying solely on pre-trained LLMs with API-level access. Moreover, FeatLLM outperforms existing fine-tuned and in-context learning approaches while maintaining lower computational costs. However, even though FeatLLM achieves state-of-the-art performance on few-shot tabular classification problems, its many iterations of rule and preprocessing function generation incur significant costs. Furthermore, during preprocessing, FeatLLM produces only binary features, which may limit expressiveness compared to the original data.
@@ 3.1 Formulation
3.2 LLM Prompting and Score Parsing
To enable Large Language Models (LLMs) to generate meaningful feature attribution scores for guiding downstream models, we developed a structured prompting methodology. This approach ensures score accuracy and relevance through three key components.
@@ 3.2.1 Task and Dataset Contextualization
We embed task and dataset details within the prompt. Task descriptions succinctly define classification objectives and outcomes, e.g., "Predict whether this patient’s breast cancer will reoccur. Yes or no?", following established methodologies [22, 24]. Feature descriptions clarify dataset attributes, e.g., "Age: The age of the patient at diagnosis.". Categorical features are one-hot encoded with explicit descriptions for each category, enabling the LLM to assign distinct attribution scores per category rather than per feature. Unlike prior approaches that use LLMs for tabular data classification, our method avoids serializing dataset examples into the prompt, thereby reducing prompt length and computational cost.
@@ 3.2.3 Score Extraction and Aggregation
A secondary LLM instance extracts numerical scores from the primary LLM’s textual output via function calling, converting semi-structured responses into a standardized list format. To enhance stability, we generate scores multiple times () and compute their mean, yielding the final LLM-based feature attribution vector ().
This methodology provides a robust framework for leveraging LLM capabilities to generate reliable feature attribution scores that effectively inform downstream predictive models.
@@ 4.1 Few-shot Learning
We evaluate our approach on ten publicly available binary classification datasets:
@@
bodyfat [43] - predicting whether an individual’s body fat percentage exceeds the mean
@@
contraceptive [53] - predicting whether an individual uses contraception
@@
indian-liver [8] - predict whether a patient has liver disease
@@
The datasets vary in size and complexity, and their additional basic information is outlined in Table 6 in the Appendix.
@@ 4.1.2 Implementation Details
We employ three distinct foundation models as our importance score estimators: Llama 3.3 70B () [16], Gemini 2.0 Flash () [44] and GPT-4o-mini () [27]. As simple downstream models, we utilized logistic regression and a two-layer multilayer perceptron (MLP) with ReLU activation and 100 hidden units. Optimization was performed using the Adam optimizer [31] with a learning rate of and no weight decay, leveraging LAAT’s internal regularization via the factor, which was set to 100. Additionally, the number of importance score estimations for LAAT was set to 5. LAAT was executed without early stopping. Data preprocessing involved one-hot encoding categorical variables and standardizing numerical features using z-score normalization. Feature descriptions were derived from dataset repository metadata or original publications, whereas task descriptions were formulated based on prior research on tabular data classification using LLMs [22, 24]. As prompt engineering is not the focus of our work, we leave the exploration of alternative prompting strategies to future research. Exact task descriptions used can be found in Table 7 in the Appendix.
For traditional machine learning models (logistic regression, MLP, random forest, XGBoost, and CatBoost), we performed hyperparameter optimization through grid search with 5-fold cross-validation, with the exception of 1-shot learning scenarios where default scikit-learn [42] hyperparameters were applied. The complete baseline hyperparameter search spaces are detailed in Table 8 in the Appendix. TabPFN and FeatLLM were implemented with their respective default parameter configurations as described in the original publications. Additionally, FeatLLM incorporated early stopping mechanisms in the 5-shot and 10-shot experimental settings. Due to the inherent complexity of generating multiple conditions and preprocessing functions in the FeatLLM method, Llama 3.3 exhibited inconsistencies in producing valid outputs. As a result, we excluded it from the FeatLLM experiments.
All models were trained using k-shot examples as the training set, while the remaining data served as the test set. To ensure statistical robustness, we conducted twenty independent experimental runs for each model. The significance of the difference in the mean values of the total scores was verified using the Wilcoxon signed-rank test with .
In the following sections, we present our experimental results and provide additional analyses on LAAT’s robustness to noise, as well as an examination of how LAAT influences the training loss landscape through the lens of the bias-variance tradeoff.
@@ 4.1.3 Results
The results of our experiments, averaged over twenty repetitions, are presented in Tables 1 (for non-LLM methods) and 2 (for LLM-based methods). As demonstrated, LAAT models consistently rank among the top-performing methods, securing the highest or shared highest scores in 28 out of 30 experiments. Notably, the highest-performing variants of LAAT were and , securing 12 and 21 top or shared top scores respectively. Furthermore, the LAAT-trained models statistically significantly outperformed their non-LAAT counterparts in 24, 28, and 27 out of 30 experiments using logistic regression with Llama 3.3 70B, Gemini 2.0 Flash, and GPT-4o-mini as scoring models, respectively. For MLP, LAAT models significantly exceeded the performance of vanilla MLP in 23, 28, and 29 experiments, respectively. These results underscore the positive impact of LLM attribution alignment on the generalization capabilities of even simple models, achieving significant improvements without the need for extensive hyperparameter tuning, which is often required for other approaches.
Dataset Shot | LR | MLP | RF | XGB | CatBoost | TabPFN | |
@@
Figure 1 illustrates the average ROC AUC performance across all datasets, highlighting the superior performance of LAAT model variants across all shot settings. While FeatLLM closely matches LAAT in the 1-shot scenario, its performance plateaus beyond this point. We hypothesize that this occurs for two reasons: first, FeatLLM generates binary features, which may lack the expressiveness of the continuous features present in the original dataset. Second, the serialized few-shot examples significantly increase the length of the input prompt, potentially reducing the effectiveness of the rule and preprocessing function generation procedure. This may occur because the expanded prompt causes the initial instruction and relevant data to become less prominent among the large number of tokens. In contrast, LAAT models demonstrate a substantial advantage in both the 5-shot and 10-shot settings, significantly outperforming all other methods. Among the scoring models evaluated, Gemini 2.0 Flash-Lite emerged as the top performer, with GPT-4o-mini following closely behind. Although Llama 3.3 70B lagged behind these two LLMs, it consistently matched or outperformed all baseline methods across all settings. In addition to outperforming FeatLLM, LAAT utilizes, on average, 79% fewer input tokens and 60% fewer output tokens, as shown in Table 3, demonstrating significant conservation of computational resources.
Method | FeatLLM | LAAT | ||||||
k-shot | 1 | 5 | 10 | N/A | ||||
Token Type | Input | Output | Input | Output | Input | Output | Input | Output |
Count | 37866 | 19730 | 51704 | 20067 | 62310 | 20282 | 10570 | 7990 |
4.1.4 Sensitivity to Importance Score Noise
We assess the robustness of the LAAT method by introducing controlled noise into the importance scores provided by the LLM. Specifically, we define a noise ratio and compute the perturbed importance scores, , as a linear interpolation between the original scores and randomly generated scores, , sampled from a uniform integer distribution in the range :
Figure 2 presents the average performance across all shot scenarios for both LR and MLP LAAT variants using the GPT-4o mini model under varying noise conditions. As expected, the performance of LAAT declines as the noise ratio increases. However, even with noise ratios as high as 0.6, LAAT variants continue to outperform their non-LAAT counterparts. These results demonstrate LAAT’s ability to enhance baseline model performance even when the LLM-derived importance scores are imperfect.
4.1.5 Exploring the Loss Landscape
To understand why LAAT consistently outperforms other methods, we analyze the training and test loss landscapes [35] of two logistic regression models: one trained with standard binary cross-entropy loss and another incorporating the attribution alignment loss with . We also visualize their respective training trajectories. Each model was trained using Adam for 200 epochs without early stopping or weight decay, with only five training samples per class, while the remaining dataset was used as the test set.
Figure 3 illustrates the results for the adult and bank datasets, while additional results for the breast-ljub and myocardial datasets are provided in the Appendix in Figure 6.
A key observation from these visualizations is that the loss landscape of the standard model exhibits a substantial discrepancy between training and test loss surfaces. This suggests high variance, where the model overfits to the limited training data, learning feature importance patterns that do not generalize well. Conversely, introducing the attribution alignment loss alters the training loss landscape, shifting it closer to the test loss landscape. This shift can be interpreted as a form of inductive bias, where the model is nudged toward an attribution structure informed by the LLM’s prior knowledge.
From a bias-variance perspective, LAAT’s regularization constrains the model’s optimization path, discouraging reliance on spurious correlations and guiding it toward more stable, generalizable parameters. Secondly, the model becomes less sensitive to variations in small datasets, lowering the variance, which is preferable in low-data scenarios. This trade-off mitigates overfitting and enhances generalization, ultimately leading to improved predictive performance, as demonstrated in previous experiments.
@@ 4.2 Learning on Biased Data
bodyfat - Exclude all individuals under the age of 50 with above-average body fat percentage, introducing a bias that overestimates body fat in elderly individuals.
@@
contraceptive - Exclude all working women not using contraceptives and all non-working women using contraceptives, creating a spurious relationship between employment and contraceptive use.
@@
indian-liver - Remove all male patients with diagnosed liver disease, biasing models toward associating liver disease more strongly with female patients.
@@ 4.2.1 Results
As in the few-shot experiment, results are presented in Tables 4 for non-LLM methods and 5 for LLM-based methods. LAAT effectively mitigates dataset biases and achieves the highest performance across all datasets, surpassing non-LLM baselines by more than 20 ROC AUC points in some cases. Although FeatLLM outperforms traditional baselines, it achieves shared top performance with LAAT on only two datasets and lags significantly behind LAAT models on the remaining datasets. The superior performance of LAAT on this challenging benchmark further highlights the benefits of attribution guidance — not only in enhancing performance in data-scarce scenarios but also in addressing subtle, hard-to-detect dataset biases that could otherwise compromise the integrity of the machine learning pipeline.
@@ 4.3 Hyperparameter Analysis
We investigate the impact of varying the regularization factor and the number of importance score estimations on the final performance of LAAT models. Specifically, we assess the performance of a logistic regression model trained using LAAT alignment with importance scores provided by GPT-4o mini. The experimental results are presented in Figures 5 and 5, corresponding to the variation and the number of importance score estimations, respectively.
@@
5 Limitations
While LAAT offers significant benefits, it has several limitations. First, the method is restricted to tabular data and supports only binary classification tasks. Additionally, LAAT requires features to be describable in natural language, limiting its applicability to datasets with anonymized or uninterpretable features. Finally, its effectiveness depends on the general knowledge embedded in LLMs, which may vary across different data domains.
@@ 6 Conclusion
Beyond its standalone effectiveness, LAAT is highly versatile and can be used in conjunction with more complex models, provided they are differentiable — for example, TabPFN. Additionally, it can be combined with feature preprocessing techniques such as FeatLLM. Given the interpretability of LLM-derived importance scores, LAAT could be incorporated into interactive, chat-based interfaces, enabling human experts to refine these scores before finalizing them. This human-in-the-loop refinement could further enhance model performance by leveraging domain expertise.
While we focus on input gradient-based attribution in this work, future research could explore alternative attribution methods to further enhance LAAT’s effectiveness. Additionally, future research could explore extending LAAT to multiclass classification problems by generating LLM-based feature attribution vectors for each class. Subsequent training could regularize local attributions based on the class of the current example. Similarly, in regression tasks, the score generation prompt of LAAT could be adapted to accommodate continuous target variables.
Furthermore, LAAT’s framework could be expanded beyond tabular data to other modalities, such as images, by integrating it into concept bottleneck models [32]. These architectures produce high-level, human-interpretable features that could serve as input to a LAAT-augmented classifier, facilitating improved feature selection and attribution in vision tasks. Such an extension could enhance both interpretability and generalization in image-based applications. We leave the exploration of these research directions to future work.
@@
We provide dataset details, task descriptions, score extraction prompt, hyperparameter search spaces, and loss landscapes used in our experiments.